AgCl has the same structure as that of NaCl .The edge length of unit cell of AgCl is found to be 555pm and the density of AgCl is 5.561g cm ^-3.find the percentage of site that are unoccupied

Grade:11

1 Answers

Arun
6 years ago
density of a crystal is mass/volumehere mass increases due to filling of impurity thus density increasespercentage of site that are unoccupiedno of agcl atom in unit cell=4density=5.561g/cm^3volume of unit cell=17*10^-23percentage unoccupied=100-percentage occupied=100-{(4*143.5/Na)/5.561}/17*10^-23= 0.24% RegardsArun(askIITians forum expert)